OpenStudy (anonymous):

what is the braking distance if travelling 60km/h?

OpenStudy (anonymous):

I didnt understand it well !

OpenStudy (dumbcow):

braking distance is proportional to speed squared...but you have to look up exact formula

OpenStudy (anonymous):

60km / h = 60x1000/3600 = 18.05m/s

OpenStudy (anonymous):

what formula am i supposed to use ??

OpenStudy (anonymous):

velocity is 18.05m/s

OpenStudy (yttrium):

Well, if @dumbcow said it correctly. I think you need to do is to find the square root of your speed. :))

OpenStudy (anonymous):

how can i do that?

OpenStudy (yttrium):

You can use your calculator, if you're allowed. :D

OpenStudy (dumbcow):

ignore reaction time in formula \[d = \frac{v^{2}}{\mu g}\]

OpenStudy (dumbcow):

oops i forgot 2 in denominator

OpenStudy (anonymous):

@Yttrium your not helpful at all

OpenStudy (anonymous):

@dumbcow your a champ

OpenStudy (anonymous):

the answer in the textbook is 13.89m however im geting a different answer

OpenStudy (dumbcow):

check your conversion....60 km/hr = 16.666 m/s

OpenStudy (anonymous):

D=(16.66)^2/2*u*g

OpenStudy (anonymous):

???

OpenStudy (dumbcow):

it could be a different "u" value...does book say anything about road surface or friction coefficient?

OpenStudy (anonymous):

no it doesnt

OpenStudy (anonymous):

something is missing from the formula

ganeshie8 (ganeshie8):

looks book wants u to use g=10

OpenStudy (dumbcow):

yeah and u =1 thats only way you get their answer

ganeshie8 (ganeshie8):

assuming road is concrete, and tyres are rubber u = 1 between Concrete & Rubber

OpenStudy (dumbcow):

the question should have specified those constants for you

OpenStudy (anonymous):

it hasnt i hate these textbooks

OpenStudy (dumbcow):

using g=10 is not very precise

OpenStudy (anonymous):

it should be 9.81 but they should of specified